asifiqbal2:

The ends of a tuning fork oscillator at a frequency of 440Hz with an amplitude of 0.50mm.determine the maximum velocity and the maximum acceleration of the ends.

SemiDefinite:

assume SHM. EG: let \(x = A \cos(\omega t ) \) \(\dot x = - \omega A \sin(\omega t ) \) \(\ddot x = - \omega^2 A \cos(\omega t ) \) \(f = 400 Hz, A = 0.50 mm\) Finally: - \(\omega = 2 \pi f\) Finish!